# Quantized polarization in a generalized Rice-Mele model at arbitrary filling

### v1

@@ IV Genralization and bulk-boundary correspondence
$U_{\rm intra}$ $=e^{i2\pi\hat{P}_{x}^{\rm intra}},\quad\hat{P}_{x}^{\rm intra}=\frac{2\pi}{L_{q}}\sum_{k=0}^{L_{q}-1}\frac{1}{q}\sum_{l=0}^{q-1}l(n_{kq+l}-\nu),$ (28)
$U_{\rm inter}$ $=e^{i2\pi\hat{P}_{x}^{\rm inter}},\quad\hat{P}_{x}^{\rm inter}=\frac{2\pi}{L_{q}}\sum_{k=0}^{L_{q}-1}k\sum_{l=0}^{q-1}(n_{kq+l}-\nu).$ (29)
@@
Finally, we remark that the existence of a gapless (single-particle) edge mode or more generally degenerate ground states under the open boundary condition could also be understood as the filling anomaly  Benalcazar et al. (2019); Tada and Oshikawa (2023). We discuss the ground state wavefunctions when $\delta=0,\Delta\gg t$ in view of the filling anomaly to have an intuitive understanding. Let us consider a simple case with $j_{0}=1$ for which $P_{x}=1/2$ is acheived when $\nu=1/3(p=1,q=3)$. There are degenerate potential minima at $j=0$ and $j=2$ (mod 3) in a unit cell and the middle site $j=1$ (mod 3) corresponds to the potential maximum. Here, we define a unit cell as the sites $\{3k,3k+1,3k+2\},k=0,1,2\cdots,L-1$ for which the bonds $\braket{3k-1,3k}$ including $\braket{L-1,0}$ connect neighboring unit cells. The ground state under the periodic boundary condition is
|$\ket{\Psi}$ $\simeq\frac{1}{\sqrt{2}}(c_{L-1}^{\dagger}+c_{0}^{\dagger})\prod_{k=0}^{L_{3}-2}\frac{1}{\sqrt{2}}(c_{3k+2}^{\dagger}+c_{3k+3}^{\dagger})\ket{0},$ (32)
where $L_{3}=L/3\in{\mathbb{Z}}$. A dimer $(c_{j}^{\dagger}+c_{j+1}^{\dagger})\ket{0}$ is formed by the hopping $t>0$ between neighboring unit cells, which is analogous to the standard SSH model with $P_{x}=1/2$ where there are dimers on inter unit cell bonds. (See Appendix D for direct calculations of the polarization in the ground state (32).) The local excitated state on the dimer bond $\braket{j,j+1}$ is $(c_{j}^{\dagger}-c_{j+1}^{\dagger})\ket{0}$ with an energy gap $2t$ which has a different inversion eigenvalue from that of the ground state. Under the open boundary condition without a hopping integral on the bond $\braket{L-1,0}$, the energy gap between the two states $(c_{L-1}^{\dagger}+c_{0}^{\dagger})\ket{0}$ and $(c_{L-1}^{\dagger}-c_{0}^{\dagger})\ket{0}$ vanishes and the state
@@ Appendix C Robustness of the charge polarization to boundary conditions
One would naively expect that the charge polarization $P_{x}={\rm arg}\bra{\Psi}U_{x}\ket{\Psi}/2\pi$ is robust to boundary conditions in an insulator. It was shown previously that the expectaion value of a local operator whose support is far way from the boundary is almost unchanged by boundary conditions in a gapped ground state based on the exponential decay of correlation functions  Watanabe (2018). However, the operator $U_{x}$ is non-local and some discussions may be necessary for the robustness of $\braket{U_{x}}$.

### v2

@@ IV Genralization and bulk-boundary correspondence
$U_{\rm intra}$ $=e^{i2\pi\hat{P}_{x}^{\rm intra}},\quad\hat{P}_{x}^{\rm intra}=\frac{1}{L_{q}}\sum_{k=0}^{L_{q}-1}\frac{1}{q}\sum_{l=0}^{q-1}l(n_{kq+l}-\nu),$ (28)
$U_{\rm inter}$ $=e^{i2\pi\hat{P}_{x}^{\rm inter}},\quad\hat{P}_{x}^{\rm inter}=\frac{1}{L_{q}}\sum_{k=0}^{L_{q}-1}k\sum_{l=0}^{q-1}(n_{kq+l}-\nu).$ (29)
@@
Finally, we remark that the existence of a gapless (single-particle) edge mode or more generally degenerate ground states under the open boundary condition could also be understood as the filling anomaly  Benalcazar et al. (2019); Tada and Oshikawa (2023). We discuss the ground state wavefunctions when $\delta=0,\Delta\gg t$ in view of the filling anomaly to have an intuitive understanding. Let us consider a simple case with $j_{0}=1$ for which $P_{x}=1/2$ is acheived when $\nu=1/3(p=1,q=3)$. There are degenerate potential minima at $j=0$ and $j=2$ (mod 3) in a unit cell and the middle site $j=1$ (mod 3) corresponds to the potential maximum. Here, we define a unit cell as the sites $\{3k,3k+1,3k+2\},k=0,1,2\cdots,L_{3}-1$ for which the bonds $\braket{3k-1,3k}$ including $\braket{L-1,0}$ connect neighboring unit cells, where $L_{3}=L/3\in{\mathbb{Z}}$. The ground state under the periodic boundary condition is
|$\ket{\Psi}$ $\simeq\frac{1}{\sqrt{2}}(c_{L-1}^{\dagger}+c_{0}^{\dagger})\prod_{k=0}^{L_{3}-2}\frac{1}{\sqrt{2}}(c_{3k+2}^{\dagger}+c_{3k+3}^{\dagger})\ket{0}.$ (32)
A dimer $(c_{j}^{\dagger}+c_{j+1}^{\dagger})\ket{0}$ is formed by the hopping $t>0$ between neighboring unit cells, which is analogous to the standard SSH model with $P_{x}=1/2$ where there are dimers on inter unit cell bonds. (See Appendix D for direct calculations of the polarization in the ground state (32).) The local excitated state on the dimer bond $\braket{j,j+1}$ is $(c_{j}^{\dagger}-c_{j+1}^{\dagger})\ket{0}$ with an energy gap $2t$ which has a different inversion eigenvalue from that of the ground state. Under the open boundary condition without a hopping integral on the bond $\braket{L-1,0}$, the energy gap between the two states $(c_{L-1}^{\dagger}+c_{0}^{\dagger})\ket{0}$ and $(c_{L-1}^{\dagger}-c_{0}^{\dagger})\ket{0}$ vanishes and the state
@@ Appendix C Robustness of the charge polarization to boundary conditions
One would naively expect that the charge polarization $P_{x}={\rm arg}\bra{\Psi}U_{x}\ket{\Psi}/2\pi$ is robust to boundary conditions in an insulator. Indeed, we have numerically checked that $P_{x}$ of the present non-interacting model does not change under the periodic boundary condition and the open boundary condition. It was shown previously that the expectaion value of a local operator whose support is far way from the boundary is almost unchanged by boundary conditions in a gapped ground state based on the exponential decay of correlation functions  Watanabe (2018). However, the operator $U_{x}$ is non-local and some discussions may be necessary for the robustness of $\braket{U_{x}}$.